Calculate the mass of oxygen that reacted with the magnesium
mass of crucible, lid and contents after reaction – mass of crucible + lid + magnesium
1
Expert's answer
2020-05-07T14:08:50-0400

Magnesium reacts with oxygen to produce magnesium oxide:

2Mg(s) + O2(g) ----> 2MgO(s)

Number of moles of Mg = mass(Mg) / Mr(Mg) = mass(Mg) / 24.3

Looking at the balanced chemical equation, 2 mol of Mg reacts with 1 mol of O2, so n mol of Mg reacts with n/2 mol of O2

mass of O2Mr × number of moles of O2 = 32.0 x n/2